A scuba tank contains 20.0 L of air at 30°C and 1.5 atm. Assuming that the temperature remains constant, what volume of air is required to fill the tank to a pressure of 40 atm? 


1
Expert's answer
2021-05-19T02:28:29-0400

"V_1=20L,P_1=1.5atm\n,\nP_2=40 atm"


using Boyle's law-


"P_1V_1=P_2V_2


\\[9pt]
\Rightarrow 20\times 1.5=V_2\times 40


\\[9pt]
\Rightarrow V_2=\dfrac{30}{40}=0.75L"